# Exploring Diverse Methods in Visual Question Answering

### v2

@@ I Introduction
In essence, the VQA task epitomizes the interdisciplinary nature of contemporary AI research [2, 3, 4, 5, 6, 7], drawing upon insights from computer vision [8, 9, 10, 11, 12], natural language processing [13, 14, 15, 16], and machine learning [17, 18, 19, 20, 21, 22, 23] to realize the ambitious goal of imbuing machines with human-like perceptual and cognitive abilities. By synthesizing information gleaned from visual stimuli with linguistic cues, VQA systems aspire to emulate the nuanced understanding and reasoning capabilities exhibited by human agents when confronted with multimodal inputs.
@@ II-A Architecture
Similar to most existing approaches to the VQA problem, we first passed the image data through a pre-trained ResNet to generate a feature vector for each image. In parallel, we also pre-processed the questions and pass them through an RNN in order to obtain feature vectors for each question. While we do not train the ResNet, we do train the RNN while training the GAN. After generating the feature vectors, we encode them together into a single vector. We attempted two approaches to combining them. Our first approach, motivated by [26, 27, 28], first passed the image and question embeddings through separate $tanh$ layers, then performed element-wise multiplication of these output vectors, and concatenated them with a vector of the element-wise attention of the outputs of the $tanh$ layers. This vector is finally passed through another $tanh$ nonlinearity. Our second approach was much simpler, utilizing separate $tanh$ layers to encode the image and question feature vectors and then combining these embeddings by element-wise multiplication.
@@ II-B Training
We attempted several variations of this. We experimented with pre-training both the Discriminator and the Generator. When pre-training the Generator we simply trained it as a softmax classifier with noise added to the inputs. This allows us to initialize the weights to values that will produce relatively good results at the start of training the full GAN [29, 30, 31, 32, 33].
We found that pre-training the Discriminator to optimality could be detrimental to the actual training process of the GAN and could worsen the updates of the Generator over time [34]. [34] shows that if the probability densities are either disjoint or lie on a low-dimensional manifold then the Discriminator can distinguish between them perfectly. This happens to be the case for the loss functions proposed by [35]. So instead of pre-training our Discriminator to optimality, we follow the suggestion in [36, 37, 38, 39, 40, 41, 42] and add noise to its inputs.
@@ IV Attention Based Mechanism
Inspired by the strength of Multimodal Compact Bilinear Pooling (MCB) at efficiently and expressively combining multimodal features [43], we use the MCB operation to replace the simple addition operation used in co-attention mechanism [44] when combining the features learned from the images and questions together, which may help to learn more information from visual part [45, 46, 47].
@@ VI Conclusion
Firstly, our experiments with GANs, especially those with full, multi-layer generator models ($GAN_{full}$), showed a marked improvement over baseline methods when pretraining is applied selectively—specifically, pretraining the generator but not the discriminator. This approach allowed the GANs to better capture and reproduce the data distribution needed for accurate VQA, although it still presents challenges in task specificity and learning the correct answers to more complex questions, aligned with previous studies [48, 49, 50, 51].

### v3

@@ I Introduction
In essence, the VQA task epitomizes the interdisciplinary nature of contemporary AI research [2], drawing upon insights from computer vision [3], natural language processing [4], and machine learning [5] to realize the ambitious goal of imbuing machines with human-like perceptual and cognitive abilities. By synthesizing information gleaned from visual stimuli with linguistic cues, VQA systems aspire to emulate the nuanced understanding and reasoning capabilities exhibited by human agents when confronted with multimodal inputs.
@@ II-A Architecture
Similar to most existing approaches to the VQA problem, we first passed the image data through a pre-trained ResNet to generate a feature vector for each image. In parallel, we also pre-processed the questions and pass them through an RNN in order to obtain feature vectors for each question. While we do not train the ResNet, we do train the RNN while training the GAN. After generating the feature vectors, we encode them together into a single vector. We attempted two approaches to combining them. Our first approach, motivated by [8, 9], first passed the image and question embeddings through separate $tanh$ layers, then performed element-wise multiplication of these output vectors, and concatenated them with a vector of the element-wise attention of the outputs of the $tanh$ layers. This vector is finally passed through another $tanh$ nonlinearity. Our second approach was much simpler, utilizing separate $tanh$ layers to encode the image and question feature vectors and then combining these embeddings by element-wise multiplication.
@@ II-B Training
We attempted several variations of this. We experimented with pre-training both the Discriminator and the Generator. When pre-training the Generator we simply trained it as a softmax classifier with noise added to the inputs. This allows us to initialize the weights to values that will produce relatively good results at the start of training the full GAN [10].
We found that pre-training the Discriminator to optimality could be detrimental to the actual training process of the GAN and could worsen the updates of the Generator over time [11]. [11] shows that if the probability densities are either disjoint or lie on a low-dimensional manifold then the Discriminator can distinguish between them perfectly. This happens to be the case for the loss functions proposed by [12]. So instead of pre-training our Discriminator to optimality, we follow the suggestion in [13] and add noise to its inputs.
@@ IV Attention Based Mechanism
Inspired by the strength of Multimodal Compact Bilinear Pooling (MCB) at efficiently and expressively combining multimodal features [14], we use the MCB operation to replace the simple addition operation used in co-attention mechanism [15] when combining the features learned from the images and questions together, which may help to learn more information from visual part [16, 17].
@@ VI Conclusion
Firstly, our experiments with GANs, especially those with full, multi-layer generator models ($GAN_{full}$), showed a marked improvement over baseline methods when pretraining is applied selectively—specifically, pretraining the generator but not the discriminator. This approach allowed the GANs to better capture and reproduce the data distribution needed for accurate VQA, although it still presents challenges in task specificity and learning the correct answers to more complex questions, aligned with previous studies [18].